OpenStudy (anonymous):

46 is what percent of 107

OpenStudy (anonymous):

107*X=46 x=46/107 x=.4299 which equates to 42.99%

OpenStudy (texaschic101):

is/of = %/100 46/107 = x/100 cross multiply 107x = 4600 x = 4600/107 x = 42.99%